# Task and Motion Planning in Hierarchical 3D Scene Graphs

### v1

@@ V Related Work
There has been recent work in the area of deriving PDDL representations for task planning from scene graphs [1], with a particular focus on using the hierarchical nature of scene graphs to sparsify the representation in order to make planning tractable. The approach used by Agia et al. [1] is only guaranteed to produce valid solutions for very specific planning domains, where only constraints between symbols with a clear “ancestor” relationships are expressible, and it is unclear how to extend this to a more general set of planning tasks. Unfortunately the scene graph abstraction may not satisfy the property of downward-refinement [2], breaking many of the assumptions in task planning. The existance of low-level geometric constraints requires going beyond task planning approaches, and into the realm of TAMP.
There has been additional work in pruning superfluous elements of scenes to accelrate TAMP. Silver et al. [19] learn to predict which symbols are relevant to a particular TAMP problem. Similarly, Srivastava et al. [20] attempt to guide TAMP from failed motion plans, much like we do in Sec III-D, however do so by adding additional goal conditions, which can complicate planning.
VI CONCLUSIONS
@@
In this work we proposed an approach for enabling and accelerating TAMP in large scene graphs. To do this, we defined characteristics of planning domains that permit the pruning of certain symbols. Then, we proposed a method for deriving a domain from a Hydra scene graph which has these characteristics, and demonstrate how we prune Places and Objects from the domain. We also prove how the plans we produce from this pruned scene graph are valid and conform to the constraints of the full planning domain. Finally, we demonstrate experimentally how our approach scales with scene graph size, goal complexity, and geometric constraints in several environments, including a scene graph built from the KITTI dataset.
@@
Acknowledgements
This work was partially funded by the ARL DCIST program, by Lincoln Laboratory’s Autonomy Al Fresco program, and by the National Defense Science and Engineering Graduate Fellowship program.

### v2

@@ 5 Related Work
There has also been recent work focused on applying structured planning domains to 3D scene graphs. Agia et al. [1] derive a PDDL representation for task planning from scene graphs, with a focus on using the hierarchical nature of scene graphs to sparsify the representation in order to make planning tractable. Their approach is only guaranteed to produce valid solutions for very specific planning domains, where only constraints between symbols with a clear “ancestor” relationships are expressible, and it is unclear how to extend this to a more general set of planning tasks. Dai et al. [7] present a method for grounding natural language commands in LTL formula, leveraging the hierarchy of the scene graphs to accelerate planning. Unfortunately the scene graph abstraction may not satisfy the property of downward-refinement [3], breaking many of the assumptions in symbolic task planning. The existence of low-level geometric constraints requires going beyond task planning approaches, and into the realm of TAMP.
There has been additional work in pruning superfluous elements of scenes to accelerate TAMP. Silver et al. [24] learn to predict which symbols are relevant to a particular TAMP problem. Khodeir et al. [18] and Vu et al. [28] both address PDDLStream’s poor scaling as the number of objects grows, with [18] proposing an algorithm that guides the search through task skeletons based on failures by the motion planner and [28] introducing a more intelligent method for instantiating streams. Meanwhile, [17] and [6] learn to guide search through predictions of relevance or feasibility. These approaches are complimentary to our own, though learning in large scene graphs suffers from issues of generalization. Outside of learning based approaches, Srivastava et al. [25] attempt to guide TAMP from failed motion plans (much like we do in Sec 3.4) by adding additional goal conditions, an approach that struggles with large initial object sets.
6 CONCLUSIONS
@@
In this work we proposed an approach for enabling and accelerating TAMP in large scene graphs. We defined characteristics of planning domains that permit the pruning of certain symbols. We then proposed a method for deriving a domain from a Hydra scene graph which has these characteristics, and demonstrate how we prune Places and Objects from the domain. We also proved how the plans we produce from this pruned scene graph are valid and conform to the constraints of the full planning domain. Finally, we demonstrated experimentally how our approach scales with scene graph size, goal complexity, and geometric constraints in several environments, including a scene graph built from the KITTI dataset and real-world execution on a Spot quadraped.
@@
Appendix
Proof of Prop 1. Consider $\pi\in\Pi_{R}$. If $\pi$ does not contain any actions parameterized by $x$, then the same plan $\pi$ is also a valid solution for $R^{\prime}$. Consider the alternative, where $\pi$ does contain an action parameterized by $x$. By Def. 1, there is another plan $\pi^{\prime}$ with equivalent motion sequence not parameterized by $x$, which is a valid solution for $R^{\prime}$.
Now that we have shown that $\Pi_{R^{\prime}}$ is not empty, we need to show that any valid plan for $R^{\prime}$ is valid for $R$. Consider plan $\pi=[a_{1},...,a_{N}]\in\Pi_{R^{\prime}}$ with corresponding state plan $\mathcal{I}_{\pi}=[\mathcal{I}_{0},...,\mathcal{I}_{N}]$. If the addition of facts $\mathcal{F}$ parameterized by $x$ make $\pi$ invalid, then there must exist a state $\mathcal{I}_{k}$ such that $\mathcal{I}_{k}\cup\mathcal{F}\notin\text{Pre}(a_{k+1})$, which means that $a_{k+1}$ is parameterized by a symbol that did not exist in $R^{\prime}$. Only a universal or existential quantifier in $\text{Pre}(a_{k+1})$ can cause $a_{k+1}$ to be parameterized by an additional symbol. Adding additional facts cannot turn an existentially quantified formula from true to false. By Definition 1, $a_{k+1}$ does not have any universal quantifiers that can be parameterized by $x$ in its precondition. Thus $\pi$ must be valid for $R$.∎
Proof of Prop 2. First note that no actions in this domain have universal quantifiers, so we only need to check Definition 1.1 to show that a symbol is redundant. Consider a place $p$ such that (not (VisitedPlace $p$)) appears in the CNF of the goal. If $p$ parameterizes moveRelaxed, then (VisitedPlace $p$) is in the effects, violating the goal. Since moveRelaxed is the only action that can be parameterized by a place, no plan can parameterize $p$ and Definition 1.1 is trivially satisfied.
Next, consider a place $p$ that does not parameterize any initial or goal facts. For any plan $\pi$ with an action parameterized by $p$, let $a_{k}^{\mathcal{P}}$ denote an action parameterized by places $\mathcal{P}$, including $p$. Plan $\pi^{\prime}$ where $a_{k}^{\mathcal{P}}$ is replaced by $a_{k}^{\mathcal{P}\setminus p}$ is also valid, since state plans $\mathcal{I}_{\pi}$ and $\mathcal{I}_{\pi^{\prime}}$ only differ by a (VisitedPlace $p$) fact, and no action preconditions or goals involve this fact. As a result the command sub-sequence corresponding to $a_{k}^{\mathcal{P}}$ is also valid for $a_{k}^{\mathcal{P}\setminus p}$. So, for any $\pi$ parameterized by $p$, we can construct $\pi^{\prime}$ that has an equivalent motion sequence but does not parameterize $p$; thus $p$ is redundant. ∎